OpenStudy (anonymous):

Solve for x show your work: x² -6x -72

OpenStudy (luigi0210):

Well since both numbers have a minus we know it has to be a (x+a)(x-b)

OpenStudy (luigi0210):

In this case the negative number is greater

OpenStudy (cwrw238):

now you need to find 2 numbers whose product is -72 which when added give -6

OpenStudy (luigi0210):

What are all the factors of 72 first of all?

OpenStudy (zzr0ck3r):

(x-3)^2 = 72+9

OpenStudy (zzr0ck3r):

yet another way^

OpenStudy (luigi0210):

Is that completing the square?

OpenStudy (cwrw238):

yes

OpenStudy (luigi0210):

@Indiadabest do you know how to finish it?

OpenStudy (anonymous):

no

OpenStudy (luigi0210):

Come on now, we pretty much numbed it down for you

OpenStudy (zzr0ck3r):

x^2 - 6x-72 = x^2 -12x+6x-72 because -12*6 = -72 and -12+6 = -6 so now factor out a x from the pair of terms on the left and a 6 from the ones on the right x(x-12)+6(x-12) factor out the x-12 (x-12)(x+6)

OpenStudy (luigi0210):

It's what zz is showing.. I never actually learned that method

OpenStudy (zzr0ck3r):

or complete the square x^2-6x-72 = 0 cut the middle term in half and then do this (x-3)^2 = 0 + 72 + 9 = 81 (x-3)^2 = 81 (x-3) = -9 and x-3 = 9 x = -12 or x = 6

OpenStudy (zzr0ck3r):

the plus 9 on the rhs is making up for the 3 on the left squared

OpenStudy (luigi0210):

And if you prefer you could graph it and see where it intersects the x-axis